Suppose an experiment is performed with ultracold neutrons having velocities of 1.05 m/s.

A) What is the wavelength of such a neutron, in nanometers?

B) What is its kinetic energy, in electronvolts?
Answer #1

Solution)

Given,

Velocity,v=1.05 m/s

Here, Momentum, P=mV=(1.675*10^-27 kg)*(1.05)

P=1.75*10^-27 kg-m/s

Now, lambda=h/P=(6.626*10^-34)/(1.75*10^-27)=3.76*10^-7 m

=376.74 nm(Ans)

==========

Part b)

Kinetic Energy, KE=P*V/2=(1.75*10^-27)*(1.05)/2=9.18*10^-28 J

=5.74*10^-9 eV (Ans)

=========